2x+y=8
y=1/2x+1/2

Answer by Cromlix(4381)

Hi there,
2x+y=8
y=1/2x+1/2
Rearrange:
2x + y = 8...........Eq(1)
-1/2x + y = 1/2......Eq(2)
Multiply Eq(2) by 4
2x + y = 8...........Eq(1)
-2x + 4y = = 2.......Eq(2)
Add Eq(1) + Eq(2)
.....5y = 10
......y = 2.
Substitute y = 2 into Eq(1)
2x + y = 8
2x + 2 = 8
2x = 8 - 2
2x = 6
x = 3
{2, 3}
Hope this helps :-)